OpenStudy (anonymous):

the equation x^2+y^2-2x+6y+3=0 is equivalent to ?.... help please medal rewarded

OpenStudy (fifciol):

you can write it down like this: \[x^2-2x+1-1+y^2+6y+9-9+3=0\] I added 1 and 9 and substracted 1and 9 so I didn't change the equation. Do you know what to do next?

OpenStudy (fifciol):

how you can write x^2-2x+1 and y^2+6y+9 ?

OpenStudy (anonymous):

wait what do you mean add and sub. 1 & 9?

OpenStudy (fifciol):

i add and substract it because i want to use the square formula

OpenStudy (fifciol):

\[x^2-2x+1=(x-1)^2\]

OpenStudy (fifciol):

and \[y^2+6x+9=(y+3)^2\]

OpenStudy (fifciol):

so our equation looks now like this: \[(x-1)^2+(y+3)^2-1-9+3=0\]

OpenStudy (fifciol):

so \[(x-1)^2+(y+3)^2=7\]

OpenStudy (fifciol):

and this is equation of a circle

OpenStudy (fifciol):

with center C=(1,-3) and radius sqrt7

OpenStudy (anonymous):

i think i got it